Please help me solve this 2 step equation: [ -2/3(k + 2/5) = 5/6 ]
Answer by TimothyLamb(4379)

-2/3(k + 2/5) = 5/6
(-2/3)k + (-2/3)(2/5) = 5/6
(-2/3)k + (-4/15) = 5/6
(-2/3)k = 5/6 + 4/15
(-2/3)k = (5*15)/(6*15) + (4*6)/(6*15)
(-2/3)k = (75)/(90) + (24)/(90)
(-2/3)k = 99/90
---
k = (99/90)(-3/2)
k = -297/180 = 1.65
k = 1.65
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php